A body of unknown temperature is placed in a refrigerator at a constant temperature of 0∘F, If after 20 minutes the temperature of the body is 40∘F, and after 40 minutes the temperature of the body is 20∘F, find the initial temperature of the body.



1
Expert's answer
2021-11-22T20:09:46-0500

"\text{The change in temperature is given by }\\
\frac{d\theta}{dt}= k(0-\theta)\\
\implies \frac{d\theta}{-\theta}=kdt\\
\implies \theta = 17-ce^{-kt}-(1)\\
\text{at t = 60 seconds and $\theta = 20$, we have that}\\
-3=ce^{-60k}-(3)\\
\text{at t = 30seconds and $\theta = 27$, we have that}\\
\text{Dividing (3) by (2), we have}\\
0.3 = e^{-30k}\\
\implies k = 0.0401\\
\text{Putting k =0.0401 in 3 we have that}\\
c = -33.33\\
\text{Next, we substitute k and c in (1), at time 0 to get our solution}\\
\therefore \theta = 17+33.33e^0=50.33^0C"